OpenStudy (anonymous):

Evaluate 3 − 24 ÷ 8 + 42

OpenStudy (anonymous):

42:)

OpenStudy (anonymous):

thankss:)

OpenStudy (anonymous):

PEMDAS says you do division first, 24/8 = 3 and 3-3=0 +42= 42(: if you need to show your work

OpenStudy (anonymous):

its not the answer it has a) 0 b)2 c)16 d)24

OpenStudy (anonymous):

umm well let me work it out kk

OpenStudy (anonymous):

it has 4 to the 2 power on 4 tho

OpenStudy (anonymous):

ohh that makes more sence

OpenStudy (anonymous):

16

OpenStudy (anonymous):

yeah 16

OpenStudy (anonymous):

bye g2g

OpenStudy (phi):

fyi, if you have it has 4 to the 2 power on 4 tho you can write that as 4^2 (4 to the 2nd power)